x P(x)

0 0.2
1 0.1
2 0.3
3 0.4
Find the expected value of the probability distribution. Round to two decimal places.

Find the standard deviation of the probability distribution. Round to two decimal places.

Answer by Boreal(15235)

add x*p(x) for each
0+0.1+0.6+1.2=1.90 expected value.
-
variance is squared dev. * prob. =(-1.9)^2*0.2+(-0.9)^2*0.1+(0.1)^2*0.3+1.1^2*0.4
=0.722+0.081+0.003+0.484
=1.29
sd=sqrt(V)=sqrt (1.29)=1.14